Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

probability

Expert replies
by jzw » Wed Mar 21, 2012 11:33 am
A box contains 100 balls, numbered from 1 to 100. If three balls are selected @ random and without replacement from the boc, what is the probability that the sum of the three numbers on the balls selected from the box will be odd?

(a) 1/4
(b) 3/8
(c) 1/2
(d) 5/8
(e) 3/4

Ok. So 50 odd and 50 even. so the probability of selecting an odd or even number are both 1/2.

In order for the sum of the three numbers to be odd...

ODD + ODD + ODD = ODD which is 1/2 * 1/2 * 1/2 * 1/2 = 1/8
ODD + ODD + EVEN = EVEN which is 1/2 * 1/2 * 1/2 * 1/2 = 1/8
ODD + EVEN + EVEN = ODD which is 1/2 * 1/2 * 1/2 * 1/2 = 1/8
EVEN + EVEN + EVEN = EVEN which is 1/2 * 1/2 * 1/2 * 1/2 = 1/8

Now, @ this point you're supposed to add these together, and get 4/8 which is 1/2, I presume because you could say, "well either the first thing will happen, OR the second, OR the third, OR the fourth." And with an OR statement we add.

Great. My question is as follows: is there a shortcut for this type of probability question that I'm unaware of or do u just have to do it out?

If the question stem been the same with the exception "if two balls" instead of three, would the probability been 1/3?
Join the discussion
Source: — Problem Solving |

by seal4913 » Wed Mar 21, 2012 12:18 pm
If u have three odds in a row without replacing its not. 5 x .5 x .5

Its. 5 x 49/99 x 48/98 plus .5 x 50/99 x 49/98

Wouldn't that make the answer a?
Join the discussion

by seal4913 » Wed Mar 21, 2012 12:59 pm
The org question states with replacement which would make .5 correct
Join the discussion

by Whitney Garner » Wed Mar 21, 2012 1:19 pm
Hi jzw -

A couple of issues with your explanation:
jzw wrote:Ok. So 50 odd and 50 even. so the probability of selecting an odd or even number are both 1/2.
That is TRUE for the first ball chosen, but this isn't the case for the second and the 3rd, because we have WITHOUT replacement.
jzw wrote: In order for the sum of the three numbers to be odd...

ODD + ODD + ODD = ODD which is 1/2 * 1/2 * 1/2 * 1/2 = 1/8
ODD + ODD + EVEN = EVEN which is 1/2 * 1/2 * 1/2 * 1/2 = 1/8
ODD + EVEN + EVEN = ODD which is 1/2 * 1/2 * 1/2 * 1/2 = 1/8
EVEN + EVEN + EVEN = EVEN which is 1/2 * 1/2 * 1/2 * 1/2 = 1/8

Now, @ this point you're supposed to add these together, and get 4/8 which is 1/2, I presume because you could say, "well either the first thing will happen, OR the second, OR the third, OR the fourth." And with an OR statement we add.
If we only want the cases that give us ODD, then we don't add ALL of the choices you listed. In fact, only 2 of the cases you have listed will work:

ODD + ODD + ODD = ODD
ODD + EVEN + EVEN = ODD

Let's start with the easier case:

(ALL ODDS)
ODD + ODD + ODD = ODD
50/100 * 49/99 * 48/98 = 1/2 * 49/99 * 24/49 = 12/99

(ONE ODD, 2 EVENS)
ODD + EVEN + EVEN = ODD
50/100 * 50/99 * 49/98 = 1/2 * 50/99 * 1/2 = 25/(2*99)

**BUT wait a moment, that is assuming you pulled the odd ball first. You could have pulled it second OR third. The probabilities won't change (I'll let you prove that to yourself), but we have to add to additional options.

EVEN + ODD + EVEN = 25/(2*99)
EVEN + EVEN + ODD = 25/(2*99)


And now we're left to add them all up:

12/99 + 3*(25/(2*99))

24/198 + 75/198 = 99/198 = 1/2

** If you're not convinced on the order, go back to your list of the 4 possible outcomes: all odd, 1 even/2 odd, 2 even/1 odd, all even. If you were to take the probabilities of every outcome, then it SHOULD = 1, not 1/2!

The reason is because of the variety of ways we could pull these:
O O O
O O E
O E O
E O O
E E O
E O E
O E E
E E E

The probabilities for these would be {12/99, 25/198, 25/198, 25/198, 25/198, 25/198, 25/198, 12/99}, find the common denominators and you will see that 24 + 6*25 + 24 = 198, so THIS gives us 1 (certainty).

jzw wrote:Great. My question is as follows: is there a shortcut for this type of probability question that I'm unaware of or do u just have to do it out?
There sure is - think about the problem this way - I'm going to take any 3 balls I pick and add them up. What are the options for the sum: odd and even. Well, does one happen more frequently than the other, let's see...

o + o + o = odd
o + o + e = even (3 cases of this depending on order)
o + e + e = odd (3 cases of this depending on order)
e + e + e = even

So, it looks like there are 8 possible outcomes, and 4 will give me an odd, the other 4 will give me an even...so half the time I'll get an even sum, 1/2 the time odd...the probability is just 1/2!

Remember that "without replacement" and "order doesn't matter" puts a monkey wrench in probability problems but we can still use logic to work it out!

Hope this helps!
:)
Whit
Whitney Garner
GMAT/GRE/EA Instructor & Anxiety/Accommodations Coach
www.whitneygarner.com

Contributor to Beat The GMAT!

Math is a lot like love - a simple idea that can easily get complicated :heart-eyes:
Join the discussion

by seal4913 » Wed Mar 21, 2012 4:49 pm
Thank you for clearing up my misunderstanding. I naturally assumed that the odds would be different when the numbers were replace and when they aren't.
Join the discussion